OpenStudy (anonymous):

Which of the following is a solution to 5(√x-7)=10? Answers: x = –3 x = –5 x = 11 x = 9

OpenStudy (akshay_budhkar):

11

OpenStudy (akshay_budhkar):

you can substitue to get your answer.

OpenStudy (angela210793):

sqrx-7=2 x-7=4 x=4+7 x=11

OpenStudy (anonymous):

i can't understand the calculation

OpenStudy (angela210793):

1)divide by 5 both sides 2) raise to the 2nd power 3) i think u get the rest

OpenStudy (anonymous):

sqartx=9 or,x=81

OpenStudy (anonymous):

is it not

OpenStudy (angela210793):

no whn u raise 2 the 2nd power u get x-7=2^2 x-7=4 x=4+7 x=11

OpenStudy (anonymous):

but there is not sqaroot(x-7) sqarootx

OpenStudy (angela210793):

no it is all under sqr...:)